A dog pen is in the shape of a right triangle. The longer leg measures 15 ft. The hypotenuse is 7ft shorter than three times the length of the shorter leg. How long is the shorter leg?
Answer by checkley77(12844)

15^2+x^2=(3x-7)^2
225+x^2=9x^2-42x+49
x^2-9x^2+42x+225-49=0
-8x^2+42x+176=0
8x^2-42x-176=0
2(4x^2-21x-88)=0
2(x-8)(4x+11)=0
x-8=0
x=8 answer for the shorter kleg.
3*8-7=24-7=17 for the hypoyenuse.
proof
15^2+8^2=17^2
225+64=289
289=289